OpenStudy (anonymous):

What is the limiting reactant when 5.6 moles of aluminum react with 6.2 moles of water? (4 points) Unbalanced equation: Al + H2O → Al2O3 + H2

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

You need to calculate how much aluminum you have and how much oxygen from the water you have.

pooja195 (pooja195):

Can you do that?

OpenStudy (anonymous):

I don't think so. Is it the same as the last problem?

pooja195 (pooja195):

yea its similar

pooja195 (pooja195):

Balanced Equation: 2Al + 3H2O → Al2O3 + 3H2 5.6 mol Al x 3 mol H2 / 2 mol Al = 8.4 mol H2 6.2 mol H2O x 3 mol H2 / 3 mol H2O = 6.2 mol H2 = Lower number, so H2O is limiting reactant

OpenStudy (anonymous):

Please explain.

OpenStudy (anonymous):

I remember having to find the low number, but I don't recall how that works.

pooja195 (pooja195):

The number of molecules on the work come from the numbers used to balance the equation, just plug them in. The lowest number is the limiting reactant also means over such as : 3 mol H2 ------------ 2 mol Al

pooja195 (pooja195):

http://openstudy.com/study#/updates/52a7d0cbe4b0d967b9d1f768 look at this its less confusing

OpenStudy (anonymous):

How is this conceived? Al: 5.6/2 = 2.8 H2O: 6.2/3 = 2.1

OpenStudy (anonymous):

6.2/3 isn't 2.1

pooja195 (pooja195):

your right the person made a mistake sorry about that :(

OpenStudy (anonymous):

However, it is still lower than Al, so it would be correct. Can you explain to me why you third H2o but half Al?

pooja195 (pooja195):

im equally confused at you are now sorry i know its H2O

OpenStudy (anonymous):

I'm questioning the fact that you divide H2o by 3, but do differently to Al.

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

I have received the correct answer. Thanks again. I have yet one last question that you may assist me with, then I shall let you alone. Will you attend?

pooja195 (pooja195):

lol yeah